

********************************************************************
UARS WEEKLY STATUS REPORT - 28 November 1997
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
Happy Thanksgiving!
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
UARS CDHF Status Week Ending Nov 28, 1997
-----------------------------------------
The CDHF has received a total of 349 CD-ROM requests as of
28 November 97. Three new CD-ROM orders were received and shipped this week.
The CDHF has successfully loaded the September 1997 DAAC data into the
DAAC Transfer Tables.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day SOLSTICE (4.52) Prod 1R-1C=0L)
1 day SUSIM (6.3) Prod (1R-1C=0L)
20 days PEM (3.32) Prod (1350R-610C=740L)
3 days MLS (4.90) Eng (43R-43C=0L)
Pending CDHF software upgrades include IDL 5.0 and OpenIngres 2.0.
***********
FOT
***********
REPORTING PERIOD
This report covers 14 November 1997 (Orbit 33757, GMT Day 318) through 21
November 1997 (Orbit 33863, GMT Day 325).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 323/2135 33832 Off, low
HALOE 323/2316 33833 Off
PEM 324/1502 33842 ZEPS on
MLS 324/2139 33847 On
PEM 325/0525 33851 Start HEPS on
PEM 325/1820 33859 Complete HEPS on
HRDI 317/0623 33732 On
The beta angle is now increasing toward a maximum angle of 78.1 degrees on
27 November 1997 (DOY 331).
Solar Stellar Pointing Platform (SSPP) special Solar tracking experiments
as requested by SOLSTICE were performed on Day 321 and Day 324. The
experiment (which has been performed earlier this year) involved commanding
the SSPP to point to a particular Right Ascension and Declination, instead
of using the Solar ephemeris to track the Sun. All of the pointing times
and angles were calculated by SOLSTICE operations and confirmed by the FOT.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.2 end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and
Battery 3 of 1.2 degrees C. Additional battery performance data for each
battery on 14 November 1997 and 21 November 1997 are provided in Appendix
D.
The clock error ranged from -10.2 to +12.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 11
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% HRDI 02
% HALOE 01
% MLS 01
% PEM 07
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
TRANSPONDER FREQUENCIES:
Transponder A: 97/301 - 2287.497086
Transponder B: 97/302 - 2287.499098
GROUND SYSTEM OPERATION
No (0) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 97/324 show the observatory data loss to be
20 hours, 41 minutes, 02.292 seconds (an increase of 2.144 seconds). This
is a 0.0383 percent data loss which equals a 99.9617 percent data capture
for the mission.
The increase of 2.144 seconds occurred on Day 322 and was due to line
problems.
The scheduled MLS ON time became a point of confusion this week when the
schedule identified operational start at the beginning of Day 325 when in
fact, MLS was not warm enough for operation until 1800 UTC of Day 325. The
scheduling inconsistency is being investigated.
FOT successfully completed an operational readiness verification of the
MSOCC playback capture and replay capability. This test was conducted to
authenticate the process as a possible backup capability during the
Building 23 power outage scheduled early 1998. Playback data in Orbit
33812 (322/1311) was downlinked to MSOCC RUPS in addition to TIPIT. The
first 10 minutes of the MSOCC recorded data was replayed through the
Application Processor (AP) at realtime rate and verified by Flight
Operations.
User Planning System (UPS) software Release 11, Build 2 was promoted to
operations this week. Prior notification for this software change was
somewhat lacking. The promotion of the software was hampered by some
configuration problems, which were eventually corrected by the UPS system
manager. In addition, the new software contains a few problems regarding
report generation, but nothing that adversely impacts the normal operations
of the FOT.
The FOT is still working out the details of participating in the IP return-
link telemetry test that the network is conducting. The FOT plans to be
participating in the test by no later than 01 December 1997.
Flight Dynamics Facility (FDF) continues the generation process for a new
Star Catalog to fix the invalid precession rate for star 48000001.
The Replacement Level Zero Processor (LZPR) was successfully elevated to
operational status on 15 November 1997.
No testing of the new CMS software release (R7B1V2) occurred this week.
Investigation of the stored command verification problem - time to complete
- associated with Applications Processor (AP) software release 10.08
continued. The cause has been identified, but rectifying the problem
continues.
OTHER SIGNIFICANT EVENTS
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 17 December
1997 at 311/132708.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
will probably be required in mid-1998.
FACILITIES
no change
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 November 1997
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
33775 320/0312 97-570 OBC/CLK Clk Adj fr 0.00 to 20.6
msec/day
33832 323/2135 97-571 HRDI "Power OFF, low "
33832 323/2135 97-572 HRDI Heater delay to 25 amp. mins
33832 323/2135 97-573 OBC Clk Adj fr 20.6 to 0.00 msec/day
33833 323/2316 97-574 HALOE Power OFF
33842 324/1502 97-575 PEM Power ON ZEPS
33847 324/2139 97-576 MLS Power ON
33849 325/0151 97-577 OBC/ Power Off HEPS 1 & 2
33852 325/0708 97-579 PEM Turned on HEPS--Low Voltage
33853 325/0855 97-580 PEM HEPS to High Voltage
33855 325/1121 97-581 PEM HEPS to Level 2
33858 325/1650 97-582 PEM HEPS to Level 4
33859 325/1820 97-583 PEM HEPS to Level 5
33861 325/2005 97-584 OBC/CLK Clk Adj fr 20.6 to 41.2 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 November 1997
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
OBSBATVEON2 Y-LO Battery 2 voltage at EON 33763-67,76-810
OBSBATVEON2 R-LO Battery 2 voltage at EON 33759-62,68-71,75-76
OBSBATVEON3 Y-LO Battery 3 voltage at EON 33759-67,76-810,21-2
OBSBATVEON3 R-LO Battery 3 voltage at EON 33759-62,68-71,75-77
OBSVLBEON Y-LO Load Bus Volts at EON 33763-67,72-74,78-
810,18,21-2
HATBAT Y-HI HALOE Bolometer area temp 33831
OBSSOC2 Y-LO Battery 2 Current SOC 33759-33863
HRIFESURVTMP Y-HI HRDI IFE Survival Temp. 33834-37
HRRADTEMP Y-HI HRDI Radiometer Temp. 33834
SOSGTEVLTG Y-HI SOLSTICE G-Tube Voltage 33759
OBSVLBEON R-LO Load Bus Volts at EON 33759-62,75-77
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 November 1997
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
33775 320/03:13 (16 Nov 97) from 00.0 to 20.6 msec/day
33832 323/21:43 (19 Nov 97) from 20.6 to 00.0 msec/day
33849 325/01:58 (21 Nov 97) from 00.0 to 20.6 msec/day
33861 325/20:14 (21 Nov 97) from 20.6 to 41.2 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 November 1997
15 Nov 1997 - GMT Day 319 Beta = 34.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.80/-3.58 +0.0/ +0.0 18.7 0.000 0.0
2 +22.4/ -33.6 +7.74/+6.07 +30.2/-16.6 24.2 1.020 24.4
3 +28.0/ -22.4 +6.44/+4.83 +29.0/-16.2 24.2 1.022 23.4
21 Nov 1997 - GMT Day 325 Beta = 60.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.94/-3.68 +0.0/ +0.0 18.7 0.000 0.0
2 +11.2/ -28.0 +5.47/+3.87 +24.1/-16.1 25.4 1.019 21.8
3 +56.0/ -11.2 +4.25/+2.80 +23.0/-15.9 25.4 1.020 20.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 November 1997
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-121 FOT 33757 318/21:37 GRD ERR SOL LD
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period